# multimotion estimation problem MOT multiobject tracking

### v3

@@ 5.3.1 Recognizing Previously Occluded Motions
$\lambda_{\rm{mc}}\left\lVert\mathbf{p}^{\check{\ell{}}_{k}C_{k}}_{C_{k}}-\mathbf{p}^{\ell{}_{k}^{\prime}C_{k}}_{C_{k}}\right\rVert+\qquad\qquad\qquad$
@@
The proportionality term, $\lambda_{\rm{mc}}$, governs how much the difference in position and the difference in velocity each contribute to whether the motions are closed. The position of the newly discovered motion, $\mathbf{p}^{\ell{}_{k}^{\prime}C_{k}}_{C_{k}}$, is calculated from the centroid of labeled points in the first frame the motion was observed as in (29). The position of the extrapolated motion, $\mathbf{p}^{\check{\ell{}}_{k}C_{k}}_{C_{k}}$, can be found from the extrapolated transform, $\mathbf{T}_{\check{\ell{}}_{k}C_{k}}$, as in (1) and (2). The body-centric velocity of the newly discovered motion, $\boldsymbol{\varpi}_{\ell{}^{\prime}_{k}}$, is calculated directly by the estimator (Section 4) and compared to the extrapolated velocity, $\boldsymbol{\varpi}_{\check{\ell{}}_{k}}$. This comparison is omitted for the pose-only estimator. This newly closed trajectory is then used to correct the extrapolated pose (Section 5.3.2).
@@ 5.3.2 Correcting Occluded States
The newly closed motion is adjusted to create a consistent, smooth trajectory. Information from both the extrapolated and newly discovered trajectories is used to correct the extrapolated states from the start of the occlusion to the time of closure, $t_{k}$. The final corrected pose, $\mathbf{T}_{{\ell{}}_{k}\ell{}_{1}}$, is calculated by correcting the extrapolated pose, $\mathbf{T}_{\check{\ell{}}_{k}\ell{}_{1}}$, to the same location as the newly discovered pose, $\mathbf{T}_{\ell{}_{k}^{\prime}\ell{}_{1}}$, using the correction transform, $\mathbf{T}_{\ell{}_{k}\check{\ell{}}_{k}}$. This correction is calculated blockwise from the translation and rotation components as in (1). The translation component of the correction transform, $\mathbf{p}^{\check{\ell{}}_{k}\ell{}_{k}}_{\ell{}_{k}}$, is the difference between the observed and extrapolated centroids, $\mathbf{p}^{\ell{}_{k}^{\prime}C_{k}}_{C_{k}}$ and $\mathbf{p}^{\check{\ell{}}_{k}C_{k}}_{C_{k}}$, such that
@@
where $\mathbf{p}^{\ell{}_{k}^{\prime}\ell{}_{k}}_{C_{k}}=\mathbf{0}$ because the newly observed centroid is taken as the position of the final corrected pose. The extrapolated camera-to-object rotation, $\mathbf{C}_{\check{\ell{}}_{k}C_{k}}$, comes blockwise from
@@ 5.3.3 Updating Closed Trajectories
The occluded states between the start of the occlusion and the corrected closure state, $\mathbf{T}_{\ell{}_{k}\ell{}_{1}}$, are then updated via interpolation (Section 5.2.2). The states after the closure, $\tau>t_{k}$, are reestimated directly from measurements relative to the corrected pose, $\mathbf{T}_{\ell{}_{k}\ell_{1}}$. The reestimate follows Section 4.2 with an estimation window starting at $t_{k}$, and the associated initial sensor-to-object transform,
@@
Parameter
OMD KITTI
Window length $\left(K\right)$
8 8
Graph neighbors $\left(k\right)$
4 4
RANSAC threshold $\left(\epsilon_{\rm{th}}\right)$
4 6
RANSAC iterations ($N_{\rm{RSAC}}$)
100 100
Outlier cost $\left(\alpha\right)$
100 100
Outlier cost $\left(\beta\right)$
5 5
Assignment proportionality $\left(\lambda_{\mathrm{sm}}\right)$
0.5 0.5
Per-label cost $\left(\mu_{\ell{}}\right)$
1000 1000
Min support size $\left(N_{\text{th}}\right)$
20 20
Min trajectory length $\left(K_{\text{th}}\right)$
3 3
Motion closure threshold $\left(\epsilon_{\text{mc}}\right)$
3 6
Motion closure prop. $\left(\lambda_{\rm{mc}}\right)$
0.25 0.25
Convergence iterations $\left(N_{\mathrm{conv}}{}\right)$
3 3
@@ 6 Evaluation
The performance of MVO is evaluated on real-world stereo data from the OMD (Judd and Gammell, 2019a) and the KITTI dataset (Geiger et al., 2012). It is evaluated both quantitatively and qualitatively and the accuracy of the different estimators is discussed. MVO is also compared to VDO-SLAM (Zhang et al., 2020b), a dynamic SLAM system that also estimates the $SE\left(3\right)$ motion of all objects in the scene, which has reported better performance than CubeSLAM (Yang and Scherer, 2019) and ClusterVO (Huang et al., 2020). Before evaluation, each estimated trajectory must first be calibrated to the ground-truth trajectory frame (Section 6.1). The accuracy of MVO and VDO-SLAM are evaluated using two metrics that give different context to their performance (Section 6.2). The OMD contains several challenging multimotion scenes with ground-truth trajectory data for each motion in the scene. The swinging_4_unconstrained segment of the OMD directly tests the ability to simultaneously segment and estimate multiple independent motions in a scene (Section 6.3). The occlusion_2_unconstrained segment exhibits significant occlusion and is used to evaluate the applicability of the motion priors for extrapolation and interpolation (Section 6.4). The KITTI dataset is a driving dataset that includes ground-truth GPS/INS data for the sensor platform, but no third-party motion data. While it does not allow for the same quantitative evaluation as the OMD, it is a commonly used benchmark for evaluating navigation algorithms and the segments Drive 0005 and Odometry 03 are used here to qualitatively evaluate performance in real-world driving scenarios (Section 6.5). The MVO parameters used in the experiments are given in Table 1. Feature detection and matching were performed using LIBVISO2 (Geiger et al., 2011) and the Gauss-Newton minimization was performed with Ceres (Agarwal et al., ) using analytical derivatives (Section 4). Note that the significant differences in sensor and depth of field between the OMD segments and the KITTI segments required slightly different values for two parameters. These parameter tunings would each be appropriate across their respective datasets.
@@ 6.1 Error Calibration
Calculating the error between an estimate and the ground truth at a time, $t_{k}$, requires calibrating the trajectories at some earlier reference time, $t_{j}<t_{k}$. This makes the calibration dependent on the chosen reference time. The calibration depends on the initial transform between the estimated and ground truth object frames,
@@
where $\mathbf{T}_{A^{\text{GT}}C}$ is the transform relating the ground-truth sensor apparatus frame from the left camera frame. The transform, $\mathbf{T}_{A^{\text{GT}}C}$, is static and provided in each dataset. The estimated object frame is assigned by each estimator and is arbitrary relative to the ground-truth object frame. The calibration at any future time is given by propagating the initial calibration forward,
@@ 6.2 Error Metrics
The accuracy of the estimation techniques is assessed using two $SE\left(3\right)$ error metrics. Global odometric error measures the estimation accuracy of the entire trajectory to a given point. Relative RMS error measures the error over consecutive time steps. Both errors provide useful context for quantifying the accuracy of the trajectory estimation. Each metric is defined using the error between the estimated motion, $\mathbf{T}_{\ell{}_{k}\ell_{j}}$, and the ground-truth motion, $\mathbf{T}_{\ell{}_{k}^{\text{GT}}\ell{}_{j}^{\text{GT}}}$, at any time $t_{k}$ and calibrated via the pose at any time $t_{j}$, given as
@@
The maximum global errors for each estimator are given in LABEL:tab:evaluation:global:swinging, LABEL:tab:evaluation:global:occlusion and LABEL:tab:evaluation:global:kitti and illustrated in Figs. 8, 9, 10, 11, 12, 13, 14 and 15. The global error on a specific data segment is contextualized by the total ground-truth path length (LABEL:tab:evaluation:global:total_path_swinging, LABEL:tab:evaluation:global:total_path_occlusion and LABEL:tab:evaluation:global:total_path_kitti) and maximum trajectory displacement (LABEL:tab:evaluation:global:swinging, LABEL:tab:evaluation:global:occlusion and LABEL:tab:evaluation:global:kitti). The total path contextualizes how far a motion travels in all directions, but does not completely describe the trajectory or the difficulty in estimating it. A repetitive motion, such as swinging blocks, unwinds accumulated error and makes the total path length less relevant for evaluating the estimator. The maximum trajectory displacement contextualizes how far a motion travels in any given direction. Unlike the total path length, this metric does not differentiate between a single traversal of a given path and multiple traversals. Together, both values are useful in understanding the global error of a given estimator on a particular data segment.
@@
This provides a single measure of the average relative error independent of when any individual error occurs. The relative error (RMSRE) is compared to the RMS ground-truth frame-to-frame motion in LABEL:tab:evaluation:relative:swinging, LABEL:tab:evaluation:relative:occlusion and LABEL:tab:evaluation:relative:kitti. The ground-truth relative motion illustrates how much each object moves between each time stamp, which can become arbitrarily small for high-frame-rate sensors, even for quickly moving objects. Small ground-truth relative motions can result in estimators having arbitrarily small relative error. Both global and relative errors are decomposed into their translational (i.e., xyz) and rotational (i.e., roll-pitch-yaw) components. All errors are reported for geocentric trajectory estimates, so a portion of the error for each motion is due to error in the camera motion estimate. VDO-SLAM third-party motion estimates were provided as calibrated relative motions. These transforms are recalibrated and combined to match the global transform representation of MVO for evaluation.
@@ 6.3 Multimotion Estimation
The swinging_4_unconstrained segment of the OMD presents complex $SE\left(3\right)$ motions without significant occlusion. It tests the ability of estimators to simultaneously segment and estimate multiple motions. It consists of four blocks that both translate and rotate in pendular motions (Fig. 5). The blocks are observed by a dynamic camera, resulting in five independent $SE\left(3\right)$ motions. A 500-frame section starting at frame 10 was used in this evaluation. Two of the blocks, Block 1 (top left) and Block 3 (bottom left), have minimal rotation and swing primarily along the ground-truth world y- and x-axis, respectively. The other two blocks, Block 2 (top right) and Block 4 (bottom right), rotate significantly about the positive ground-truth body z-axis. Block 4 rotates counter-clockwise and Block 2 rotates clockwise while also swinging, which creates a complex motion.
@@ 6.5 Urban Driving
The KITTI Drive 0005 segment (Fig. 7) is a 154-frame sequence that contains two independent motions, a cyclist and van, observed by a car-mounted camera as they travel around two consecutive turns. The van is temporarily occluded multiple times as it rounds a turn, is obscured by the cyclist, or is overexposed by sunlight. These all contribute to a challenging tracking and estimation problem that is compounded by the significant distance between the van and the camera. The KITTI Odometry 03 segment, is a longer 800-frame sequence that only contains one independent motion, another car, observed for part of the segment. This segment is useful for evaluating egomotion estimation over a significant distance but is not not particularly interesting for multimotion estimation.
@@ 7 Discussion
Qualitative and quantitative evaluations show that MVO consistently estimates ego- and third-party motions, even in the presence of temporary occlusions. MVO demonstrates good estimation and tracking accuracy on the evaluated toy problems and a real-world driving scenario without relying on appearance-based detectors or preprocessing segmentation stages. MVO uses sparse feature tracking which does not require prior knowledge about object appearance but has known limitations (Section 7.1). The distribution of features on each object affects the estimation of its motion, and small or distant objects can be difficult to segment and estimate. Similarly, MVO demonstrates that motion-based tracking through occlusions is a reliable way to address the MEP without making appearance-based assumptions, but accurately estimating the rotation of a previously occluded object without modeling its shape is challenging. Some of the limitations of MVO are inherent to the estimators it employs, each bringing its own advantages and disadvantages (Section 7.2). Despite its limitations, MVO outperforms VDO-SLAM in both egomotion and third-party trajectory estimation accuracy (Section 7.3).
@@ 7.1 Limitations of Sparse Approaches
MVO is a sparse, feature-based approach and therefore its estimation accuracy depends on the distribution of features on each dynamic object. This distribution affects the nature of the estimated motion, the observed shape of the object, and the motion extrapolation during occlusions. Features are only observed on object surfaces facing the sensor, which makes it difficult to estimate certain types of motions, such as translations along the optical axis and rotations not around the optical axis. For example, Block 2 and Block 4 in the swinging_4_unconstrained segment (Figs. 8, 9 and 10) both rotate significantly around the z-axis, leading to larger rotational errors. Not only are these rotations outside the image plane, but it is also difficult to differentiate between an object rotating in place and orbiting about an offset axis. These challenges could be overcome by using longer estimation windows or incorporating 3D fusion techniques (Rünz and Agapito, 2017). Despite this, the yaw errors for both Block 2 and Block 4 are less than $7.5\%$ of their total yaw rotation for all MVO estimators. The observable shape (and centroid) of an object also changes over time as it and the sensor move. This shifts the observable features on the object and the geocentric trajectory estimate, which degrades the interpolated estimates. This is often most evident before and after an occlusion, as seen in the estimates for the block tower in the occlusion_2_unconstrained segment (Figs. 10(b), 11(b) and 12(b)). The block tower often extends outside the camera frustum while stopped (i.e., occluded for motion-based tracking), which leads to a large discrete jump in its observed centroid when it is tracked again. These discrete jumps create interpolation errors during motion closure and lead to considerably higher global errors than would otherwise be expected given the small relative error (LABEL:tab:evaluation:global:occlusion and LABEL:tab:evaluation:relative:occlusion) The transition into occlusion itself presents estimation challenges. As the object becomes more occluded, the number of tracked features decreases and the quality of the segmentation and estimation degrades. This is evident in the egomotion and block tower estimates for the occlusion_2_unconstrained segment, which are occasionally corrupted when the the block tower slows to a stop and an increasing number of its features are incorrectly classified as part of the background (Figs. 11, 12 and 13). The transition into occlusion also makes extrapolation harder, because the erroneous estimates before an occlusion are extrapolated forward and make motion closure more difficult. This is evident in the swinging block estimates for the occlusion_2_unconstrained segment, which are often poorly extrapolated (Figs. 10(c), 11(c) and 12(c)). Objects rarely become occluded or unoccluded instantaneously and performance in the transitional periods could be improved by predicting occlusions (Mitzel et al., 2010) or more sophisticated centroid calculation (Setterfield et al., 2018).

### v4

@@ 5.3.1 Recognizing Previously Occluded Motions
$\lambda_{\rm{mc}}\left\lVert\mathbf{p}^{\check{\ell{}}_{k}C_{k}}_{C_{k}}-\mathbf{p}^{\ell{}_{k}^{\prime}C_{k}}_{C_{k}}\right\rVert+$
@@
The proportionality term, $\lambda_{\rm{mc}}$, governs how much the difference in position and the difference in velocity each contribute to whether the motions are closed.
The position of the newly discovered motion, $\mathbf{p}^{\ell{}_{k}^{\prime}C_{k}}_{C_{k}}$, is calculated from the centroid of labeled points in the first frame the motion was observed as in (21). The position of the extrapolated motion, $\mathbf{p}^{\check{\ell{}}_{k}C_{k}}_{C_{k}}$, can be found from the extrapolated transform, $\mathbf{T}_{\check{\ell{}}_{k}C_{k}}$, as in (1) and (2).
The body-centric velocity of the newly discovered motion, $\boldsymbol{\varpi}_{\ell{}^{\prime}_{k}}$, is calculated directly by the estimator (Section 4) and compared to the extrapolated velocity, $\boldsymbol{\varpi}_{\check{\ell{}}_{k}}$. This comparison is omitted for the pose-only estimator.
This newly closed trajectory is then used to correct the extrapolated pose (Section 5.3.2).
@@ 5.3.2 Correcting Occluded States
The newly closed motion is adjusted to create a consistent, smooth trajectory. Information from both the extrapolated and newly discovered trajectories is used to correct the extrapolated states from the start of the occlusion to the time of closure, $t_{k}$.
The final corrected pose, $\mathbf{T}_{{\ell{}}_{k}\ell{}_{1}}$, is calculated by correcting the extrapolated pose, $\mathbf{T}_{\check{\ell{}}_{k}\ell{}_{1}}$, to the same location as the newly discovered pose, $\mathbf{T}_{\ell{}_{k}^{\prime}\ell{}_{1}}$, using the correction transform, $\mathbf{T}_{\ell{}_{k}\check{\ell{}}_{k}}$. This correction is calculated blockwise from the translation and rotation components as in (1).
The translation component of the correction transform, $\mathbf{p}^{\check{\ell{}}_{k}\ell{}_{k}}_{\ell{}_{k}}$, is the difference between the observed and extrapolated centroids, $\mathbf{p}^{\ell{}_{k}^{\prime}C_{k}}_{C_{k}}$ and $\mathbf{p}^{\check{\ell{}}_{k}C_{k}}_{C_{k}}$, such that
@@
where $\mathbf{p}^{\ell{}_{k}^{\prime}\ell{}_{k}}_{C_{k}}=\mathbf{0}$ because the newly observed centroid is taken as the position of the final corrected pose.
The extrapolated camera-to-object rotation, $\mathbf{C}_{\check{\ell{}}_{k}C_{k}}$, comes blockwise from
@@
The rotation component of the correction transform, $\mathbf{C}_{\ell{}_{k}\check{\ell{}}_{k}}$, is not estimated and therefore taken as identity, which makes the rotation of the final corrected pose, $\mathbf{C}_{{\ell{}}_{k}{\ell{}}_{1}}$, equivalent to the rotation of extrapolated pose, $\mathbf{C}_{\check{\ell{}}_{k}{\ell{}}_{1}}$. Appearance-based methods could instead be used to estimate absolute object orientation after occlusion.
@@ 5.3.3 Updating Closed Trajectories
The occluded states between the start of the occlusion and the corrected closure state, $\mathbf{T}_{\ell{}_{k}\ell{}_{1}}$, are then updated via interpolation (Section 5.2.2).
The states after the closure, $\tau>t_{k}$, are reestimated directly from measurements relative to the corrected pose, $\mathbf{T}_{\ell{}_{k}\ell_{1}}$. The reestimate follows Section 4.2 with an estimation window starting at $t_{k}$, and the associated initial sensor-to-object transform,
@@
Parameter OMD KITTI
Window length $\left(K\right)$ 8 8
Graph neighbors $\left(k\right)$ 4 4
RANSAC threshold $\left(\epsilon_{\rm{th}}\right)$ 4 6
RANSAC iterations ($N_{\rm{RSAC}}$) 100 100
Outlier cost $\left(\alpha\right)$ 100 100
Outlier cost $\left(\beta\right)$ 5 5
Assignment proportionality $\left(\lambda_{\mathrm{sm}}\right)$ 0.5 0.5
Per-label cost $\left(\mu_{\ell{}}\right)$ 1000 1000
Min support size $\left(N_{\text{th}}\right)$ 20 20
Min trajectory length $\left(K_{\text{th}}\right)$ 3 3
Motion closure threshold $\left(\epsilon_{\text{mc}}\right)$ 3 6
Motion closure prop. $\left(\lambda_{\rm{mc}}\right)$ 0.25 0.25
Convergence iterations $\left(N_{\mathrm{conv}}{}\right)$ 3 3
@@ 6 Evaluation
The performance of MVO is evaluated on real-world stereo data from the OMD (Judd and Gammell, 2019a) and the KITTI dataset (Geiger et al., 2012). It is evaluated both quantitatively and qualitatively and the accuracy of the different estimators is discussed. MVO is also compared to VDO-SLAM (Zhang et al., 2020b), a dynamic SLAM system that also estimates the $SE\left(3\right)$ motion of all objects in the scene, which has reported better performance than CubeSLAM (Yang and Scherer, 2019) and ClusterVO (Huang et al., 2020).
Before evaluation, each estimated trajectory must first be calibrated to the ground-truth trajectory frame (Section 6.1). The accuracy of MVO and VDO-SLAM are evaluated using two metrics that give different context to their performance (Section 6.2).
The OMD contains several challenging multimotion scenes with ground-truth trajectory data for each motion in the scene. The swinging_4_unconstrained segment of the OMD directly tests the ability to simultaneously segment and estimate multiple independent motions in a scene (Section 6.3). The occlusion_2_unconstrained segment exhibits significant occlusion and is used to evaluate the applicability of the motion priors for extrapolation and interpolation (Section 6.4).
The KITTI dataset is a driving dataset that includes ground-truth GPS/INS data for the sensor platform, but no third-party motion data. While it does not allow for the same quantitative evaluation as the OMD, it is a commonly used benchmark for evaluating navigation algorithms and the segments Drive 0005 and Odometry 03 are used here to qualitatively evaluate performance in real-world driving scenarios (Section 6.5).
The MVO parameters used in the experiments are given in Table 1. Feature detection and matching were performed using LIBVISO2 (Geiger et al., 2011) and the Gauss-Newton minimization was performed with Ceres (Agarwal et al., ) using analytical derivatives (Section 4). Note that the significant differences in sensor and depth of field between the OMD segments and the KITTI segments required slightly different values for two parameters. These parameter tunings would each be appropriate across their respective datasets.
@@ 6.1 Error Calibration
Calculating the error between an estimate and the ground truth at a time, $t_{k}$, requires calibrating the trajectories at some earlier reference time, $t_{j}<t_{k}$. This makes the calibration dependent on the chosen reference time.
The calibration depends on the initial transform between the estimated and ground truth object frames,
@@
where $\mathbf{T}_{A^{\text{GT}}C}$ is the transform relating the ground-truth sensor apparatus frame from the left camera frame. The transform, $\mathbf{T}_{A^{\text{GT}}C}$, is static and provided in each dataset. The estimated object frame is assigned by each estimator and is arbitrary relative to the ground-truth object frame.
The calibration at any future time is given by propagating the initial calibration forward,
@@ 6.2 Error Metrics
The accuracy of the estimation techniques is assessed using two $SE\left(3\right)$ error metrics. Global odometric error measures the estimation accuracy of the entire trajectory to a given point. Relative RMS error measures the error over consecutive time steps. Both errors provide useful context for quantifying the accuracy of the trajectory estimation.
Each metric is defined using the error between the estimated motion, $\mathbf{T}_{\ell{}_{k}\ell_{j}}$, and the ground-truth motion, $\mathbf{T}_{\ell{}_{k}^{\text{GT}}\ell{}_{j}^{\text{GT}}}$, at any time $t_{k}$ and calibrated via the pose at any time $t_{j}$, given as
@@
The maximum global errors for each estimator are given in LABEL:tab:evaluation:global:swinging, LABEL:tab:evaluation:global:occlusion and LABEL:tab:evaluation:global:kitti and illustrated in Figs. 8, 9, 10, 11, 12, 13, 14 and 15.
The global error on a specific data segment is contextualized by the total ground-truth path length (LABEL:tab:evaluation:global:total_path_swinging, LABEL:tab:evaluation:global:total_path_occlusion and LABEL:tab:evaluation:global:total_path_kitti) and maximum trajectory displacement (LABEL:tab:evaluation:global:swinging, LABEL:tab:evaluation:global:occlusion and LABEL:tab:evaluation:global:kitti). The total path contextualizes how far a motion travels in all directions, but does not completely describe the trajectory or the difficulty in estimating it. A repetitive motion, such as swinging blocks, unwinds accumulated error and makes the total path length less relevant for evaluating the estimator. The maximum trajectory displacement contextualizes how far a motion travels in any given direction. Unlike the total path length, this metric does not differentiate between a single traversal of a given path and multiple traversals. Together, both values are useful in understanding the global error of a given estimator on a particular data segment.
@@
This provides a single measure of the average relative error independent of when any individual error occurs.
The relative error (RMSRE) is compared to the RMS ground-truth frame-to-frame motion in LABEL:tab:evaluation:relative:swinging, LABEL:tab:evaluation:relative:occlusion and LABEL:tab:evaluation:relative:kitti. The ground-truth relative motion illustrates how much each object moves between each time stamp, which can become arbitrarily small for high-frame-rate sensors, even for quickly moving objects. Small ground-truth relative motions can result in estimators having arbitrarily small relative error.
Both global and relative errors are decomposed into their translational (i.e., xyz) and rotational (i.e., roll-pitch-yaw) components. All errors are reported for geocentric trajectory estimates, so a portion of the error for each motion is due to error in the camera motion estimate.
VDO-SLAM third-party motion estimates were provided as calibrated relative motions. These transforms are recalibrated and combined to match the global transform representation of MVO for evaluation.
@@ 6.3 Multimotion Estimation
The swinging_4_unconstrained segment of the OMD presents complex $SE\left(3\right)$ motions without significant occlusion. It tests the ability of estimators to simultaneously segment and estimate multiple motions. It consists of four blocks that both translate and rotate in pendular motions (Fig. 5). The blocks are observed by a dynamic camera, resulting in five independent $SE\left(3\right)$ motions. A 500-frame section starting at frame 10 was used in this evaluation.
Two of the blocks, Block 1 (top left) and Block 3 (bottom left), have minimal rotation and swing primarily along the ground-truth world y- and x-axis, respectively. The other two blocks, Block 2 (top right) and Block 4 (bottom right), rotate significantly about the positive ground-truth body z-axis. Block 4 rotates counter-clockwise and Block 2 rotates clockwise while also swinging, which creates a complex motion.
@@ 6.5 Urban Driving
The KITTI Drive 0005 segment (Fig. 7) is a 154-frame sequence that contains two independent motions, a cyclist and van, observed by a car-mounted camera as they travel around two consecutive turns. The van is temporarily occluded multiple times as it rounds a turn, is obscured by the cyclist, or is overexposed by sunlight. These all contribute to a challenging tracking and estimation problem that is compounded by the significant distance between the van and the camera.
The KITTI Odometry 03 segment, is a longer 800-frame sequence that only contains one independent motion, another car, observed for part of the segment. This segment is useful for evaluating egomotion estimation over a significant distance but is not not particularly interesting for multimotion estimation.
@@ 7 Discussion
Qualitative and quantitative evaluations show that MVO consistently estimates ego- and third-party motions, even in the presence of temporary occlusions. MVO demonstrates good estimation and tracking accuracy on the evaluated toy problems and a real-world driving scenario without relying on appearance-based detectors or preprocessing segmentation stages.
MVO uses sparse feature tracking which does not require prior knowledge about object appearance but has known limitations (Section 7.1). The distribution of features on each object affects the estimation of its motion, and small or distant objects can be difficult to segment and estimate. Similarly, MVO demonstrates that motion-based tracking through occlusions is a reliable way to address the MEP without making appearance-based assumptions, but accurately estimating the rotation of a previously occluded object without modeling its shape is challenging. Some of the limitations of MVO are inherent to the estimators it employs, each bringing its own advantages and disadvantages (Section 7.2). Despite its limitations, MVO outperforms VDO-SLAM in both egomotion and third-party trajectory estimation accuracy (Section 7.3).
@@ 7.1 Limitations of Sparse Approaches
MVO is a sparse, feature-based approach and therefore its estimation accuracy depends on the distribution of features on each dynamic object. This distribution affects the nature of the estimated motion, the observed shape of the object, and the motion extrapolation during occlusions.
Features are only observed on object surfaces facing the sensor, which makes it difficult to estimate certain types of motions, such as translations along the optical axis and rotations not around the optical axis. For example, Block 2 and Block 4 in the swinging_4_unconstrained segment (Figs. 8, 9 and 10) both rotate significantly around the z-axis, leading to larger rotational errors. Not only are these rotations outside the image plane, but it is also difficult to differentiate between an object rotating in place and orbiting about an offset axis. These challenges could be overcome by using longer estimation windows or incorporating 3D fusion techniques (Rünz and Agapito, 2017). Despite this, the yaw errors for both Block 2 and Block 4 are less than $7.5\%$ of their total yaw rotation for all MVO estimators.
@@
The observable shape (and centroid) of an object also changes over time as it and the sensor move. This shifts the observable features on the object and the geocentric trajectory estimate, which degrades the interpolated estimates. This is often most evident before and after an occlusion, as seen in the estimates for the block tower in the occlusion_2_unconstrained segment (Figs. 11(b), 12(b) and 13(b)). The block tower often extends outside the camera frustum while stopped (i.e., occluded for motion-based tracking), which leads to a large discrete jump in its observed centroid when it is tracked again. These discrete jumps create interpolation errors during motion closure and lead to considerably higher global errors than would otherwise be expected given the small relative error (LABEL:tab:evaluation:global:occlusion and LABEL:tab:evaluation:relative:occlusion)
The transition into occlusion itself presents estimation challenges. As the object becomes more occluded, the number of tracked features decreases and the quality of the segmentation and estimation degrades. This is evident in the egomotion and block tower estimates for the occlusion_2_unconstrained segment, which are occasionally corrupted when the the block tower slows to a stop and an increasing number of its features are incorrectly classified as part of the background (Figs. 11, 12 and 13).
The transition into occlusion also makes extrapolation harder, because the erroneous estimates before an occlusion are extrapolated forward and make motion closure more difficult. This is evident in the swinging block estimates for the occlusion_2_unconstrained segment, which are often poorly extrapolated (Figs. 11(c), 12(c) and 13(c)). Objects rarely become occluded or unoccluded instantaneously and performance in the transitional periods could be improved by predicting occlusions (Mitzel et al., 2010) or more sophisticated centroid calculation (Setterfield et al., 2018).